# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 3. $(C,\alpha)$-good property
Let $\delta\in(0,1]$. A family $\mathscr{F}$ of real functions defined on $[T_{0},\infty)$ is called $\delta$-good if there exists a constant $M(\delta)$ depending only on $\delta$ such that
@@
for all $f\in\mathscr{F}$ and all bounded sub-intervals $I_{\delta}\subset I\subseteq[T_{0},\infty]$ satisfying $|I_{\delta}|=\delta|I|$.
Such inequalities are well-known in the literature as Remez-type inequalities; see [Remez_ineq]. We have the following theorem, which will imply Theorem 1.14.
@@
Let $\mathcal{V}$ be a finite-dimensional vector space of real functions defined on $[0,\infty)$ which are definable in a polynomially bounded o-minimal structure. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\smallsetminus\{0\}$,
$\lim_{t\to\infty}f(t)\neq 0.$
Then, there exists $T_{0}>0$ such that for all $\delta\in(0,1]$ and all $f\in\mathscr{F}\smallsetminus\{0\}$ it holds that $f$ restricted to $[T_{0},\infty)$ is $\delta$-good.
@@
We first note the following Corollary from Theorem 3.2.
@@
Let $\mathcal{V}$ be a finite dimensional vector space of real functions on $[0,\infty)$ which are definable in a polynomially bounded o-minimal structure. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\smallsetminus\{0\}$,
@@
Let $T_{0}\geq 1$ such that the outcome of Theorem 3.2 holds. Then, there exists $\mathrm{m},r>0$ such that for all $x\geq 1$, $f\in\mathscr{F}\smallsetminus\{0\}$ and $I^{\prime}\subseteq I\subseteq[T_{0},\infty)$ such that $\frac{|I|}{|I^{\prime}|}\leq x$ it holds that
(3.2) $\|f\|_{I}\leq\mathrm{m}x^{r}\|f\|_{I^{\prime}}.$
@@
Consider the definable set:
$S:=\{(x,y):$ $x,y>0\text{, such that }\frac{\|f\|_{I}}{\|f\|_{I^{\prime}}}\leq y,\,\forall 0\neq f\in\mathscr{F},$
$I^{\prime}\subseteq I\subseteq[T_{0},\infty),\,\frac{|I|}{|I^{\prime}|}\leq x\}.$
By Theorem 3.2, the projection of $S$ to the first coordinate includes $[1,\infty).$ Then, by the definable choice theorem (see [Geom_cat_and_o_min_str, Section 4]), there exists a function $\phi:[1,\infty)\to\mathbb{R}^{2}$ definable in the same polynomially bounded o-minimal structure, such that
@@
Namely for all $x>1$, $f\in\mathscr{F}\smallsetminus\{0\}$ and $I^{\prime}\subseteq I\subseteq[T_{0},\infty)$ such that $\frac{|I|}{|I^{\prime}|}\leq x$ it holds that
@@
Since $\phi$ is polynomially bounded, there exists $r>0$ and $T_{1}\geq 1$ such that $\frac{\phi(x)}{x^{r}}\leq c_{0}$ for all $x\geq T_{1}$, for some $c_{0}>0$. Finally, since $\frac{\phi(x)}{x^{r}}$ is bounded in $[1,T_{1}]$, the result follows. ∎
We are now ready to prove the $(C,\alpha)$-good property. Our proof is based on [Klein_Marg_flows_and_dioph, proof of Proposition 3.2].
Proof of Theorem 1.14.
By o-minimality, the number of connected components of the family of definable (in parameters) sub-level sets
$\{t\in I:|f(t)|\leq\epsilon\},~{}\epsilon>0,~{}I\subset[T_{0},\infty),~{}f\in\mathscr{F}\smallsetminus\{0\},$
is bounded uniformly, say by $K$. Now fix $I\subset[T_{0},\infty),~{}\epsilon>0$ and an $f\in\mathscr{F}\smallsetminus\{0\}$. Let
@@
$L\leq K|I^{\prime}|,$
where
$L:=|\{t\in I:|f(t)|\leq\epsilon\}|.$
The latter inequality implies,
$\frac{|I|}{|I^{\prime}|}\leq\frac{|I|}{L/K}.$
By Corollary 3.3, we get
@@
The space of polynomials of bounded degrees is the basic example of $\delta$-good functions. We provide a proof for the sake of completeness.
@@
Fix $n\in\mathbb{N}$, and consider $\mathscr{F}:=\text{Span}_{\mathbb{R}}\{1,x,...,x^{n}\}\smallsetminus\{0\}$. Then $\mathscr{F}$ is $\delta$-good on $\mathbb{R}$. More precisely:
(3.3) $\frac{\|c_{0}+c_{1}x+\cdots+c_{n}x^{n}\|_{I}}{\|c_{0}+c_{1}x+\cdots+c_{n}x^{n}\|_{I_{\delta}}}\leq(n+1)\frac{n^{n}}{\delta^{n}},$
for all $c_{0},c_{1},...,c_{n}\in\mathbb{R}$ not all zero, interval $I\subset\mathbb{R}$ and sub-interval $I_{\delta}\subset I$ satisfying $|I_{\delta}|=\delta|I|>0.$
@@
Let $f(x):=c_{0}+c_{1}x+\cdots+c_{n}x^{n}$ and $I=[a,a+T]$. Let $a\leq t_{0}\leq a+T-\delta T$, and denote $t_{i}:=t_{0}+\frac{i\delta T}{n}$ for $i=0,1,...,n$, $a\leq t_{0}<t_{1}<\cdots<t_{n}\leq a+T$, and let $I_{\delta}:=[t_{0},t_{n}]$. Then by polynomial interpolation of $f$ at points $t_{0},t_{1},...,t_{n}$, we have
(3.4) $f(t)=\sum_{j=0}^{n}\frac{\prod_{i\neq j}(t-t_{i})}{\prod_{i\neq j}(t_{j}-t_{i})}f(t_{j}).$
Now,
$\left|\frac{\prod_{i\neq j}(t-t_{i})}{\prod_{i\neq j}(t_{j}-t_{i})}\right|\leq\frac{n^{n}}{\delta^{n}},~{}\forall t\in I,$
and thus, by triangle inequality,
$\left|f(t)\right|\leq(n+1)\frac{n^{n}}{\delta^{n}}\max{\left|f(t_{i})\right|}\leq(n+1)\frac{n^{n}}{\delta^{n}}\|f\|_{I_{\delta}},~{}\forall t\in I.$
We will need the following facts:
  • For any definable function $f:[0,\infty)\to\mathbb{R}$, there exits $t_{0}\geq 0$ such that $f$ differentiable on $[t_{0},\infty)$ and $f^{\prime}$ is continuous and definable on $[t_{0},\infty)$ (see [Geom_cat_and_o_min_str, Cell decompsition, Section 4]).
  • For a definable function $f:[0,\infty)\to\mathbb{R}$ in a polynomially bounded o-minimal structure with $f(t)\sim t^{r}$ where $r\neq 0$, we have $f^{\prime}(t)\sim rt^{r-1},\text{ as }t\to\infty$, see [Miller94, Proposition 3.1]. In particular, if $\deg(f)\neq 0,$ then $\deg(f^{\prime})=\deg(f)-1$.
@@
For $i=0,1,\ldots,N$, let $F_{i}:[0,\infty)\to\mathbb{R}$ be a function definable in a polynomially bounded o-minimal structure such that $F_{i}$ is not eventually the zero function. Suppose that the degrees $\deg(F_{0}),\deg(F_{1}),\ldots,\deg(F_{N})$ are all distinct. Then, there exists a $T_{0}>0$ such that for any interval $I\subset[T_{0},\infty)$ it holds that
$\|c_{0}F_{0}+c_{1}F_{1}+...+c_{N}F_{N}\|_{I}>0$
for all $(c_{0},c_{1},...,c_{N})\neq 0$.
@@
By o-minimality, the definable set $F_{0}^{-1}(0)$ has finitely many connected components. Since $F_{0}$ is not eventually the zero function, for some $T_{0}>0$, $F_{0}(t)\neq 0$ for all $t\geq T_{0}$. Thus for all $I\subset[T_{0},\infty)$ we have
$\left\|c_{0}F_{0}+c_{1}F_{1}+...+c_{N}F_{N}\right\|_{I}>0\iff\left\|c_{0}+c_{1}\frac{F_{1}}{F_{0}}+...+c_{N}\frac{F_{N}}{F_{0}}\right\|_{I}>0.$
Clearly, since the degrees of $F_{0},F_{1},...,F_{N}$ are distinct, the degrees of $1,F_{1}/F_{0},F_{2}/F_{0},\cdots F_{N}/F_{0}$ are distinct. Thus, it is sufficient to prove the statement under the assumption that $F_{0}(t)\equiv 1$.
We now prove the statement for $N=1$, and then argue by induction. Since $\deg(F_{1})\neq\deg(F_{0})=0$, we let $T_{0}>0$ be large enough such that $F_{1}^{\prime}(t)\neq 0$ for all $t\geq T_{0}$. Assume that for an interval $I\subset[T_{0},\infty)$ we have $\|c_{0}+c_{1}F_{1}\|_{I}=0$. Since $c_{0}+c_{1}F_{1}(t)=0$ for all $t\in I$, we get that
$\frac{d}{dt}\left(c_{0}+c_{1}F_{1}(t)\right)=c_{1}F_{1}^{\prime}(t)=0,~{}\forall t\in I,$
and as $F_{1}^{\prime}(t)\neq 0$ for all $t\in[T_{0},\infty)$, it follows that $c_{1}=0$. As a consequence, $c_{0}=0$.
Now let $N\geq 2$, and let $T_{1}$ such that the functions $F_{1},F_{2},...,F_{n}$ are differentiable in $[T_{1},\infty)$. Note that since the degrees of the functions $1,F_{1},F_{2},...$ $,F_{n}$ are distinct, we have that $\deg(F_{i})\neq 0$ for all $1\leq i\leq N$. In particular, $\deg(F_{i}^{\prime})=\deg(F_{i})-1$, and so the degrees of $F_{1}^{\prime},F_{2}^{\prime},\cdots,F_{N}^{\prime}$ are all distinct.
Now, if
$\|c_{0}+c_{1}F_{1}+...+c_{N}F_{N}\|_{I}=0,$
then
$\|c_{1}F_{1}^{\prime}+...+c_{N}F_{N}^{\prime}\|_{I}=0.$
The claim now follows by induction. ∎
Note that if $F_{0}(t),F_{1}(t),...,F_{N}(t)$ are definable functions, then they are continuously differentiable $(N+1)$ times for all $t\geq T$ for some $T$. Hence, the Wronskian matrix $W(F_{0},...,F_{N})(t)$ which is the $(N+1)\times(N+1)$ matrix whose $k$-th row is $(F^{(k)}_{0}(t),F_{1}^{(k)}(t),\ldots,F_{N}^{(k)}(t))$ for $0\leq k\leq N$ is well defined for all $t\geq T$.
@@
Suppose that $F_{i}:[0,\infty)\to\mathbb{R},~{}i=0,1,...,N$, are functions definable in a polynomially bounded o-minimal structure such that the degrees $\deg(F_{i})$, for $0\leq i\leq N$ are distinct. Then, there exists a $T_{0}>0$ such that the functions $F_{0},...,F_{N}$ are $(N+1)$-time continuously differentiable and $W(F_{0},...,F_{N})(t)$ is non-singular for all $t\geq T_{0}$.
Proof.
Since $\det(W(F_{0},...,F_{N})(t))$ is a definable function, then either
$|\det(W(F_{0},...,F_{N})(t))|>0,$
for all large $t$ or $\det(W(F_{0},...,F_{N})(t))=0$ for all large $t$. Suppose for contradiction the latter case, say $\det(W(F_{0},...,F_{N})(t))=0,\forall t\geq T_{1}$. Then, by a classical result of Bôcher [Bochner_certain_cases_Wronksian_implies_dep], we get that $F_{0},F_{1},...,F_{N}$ are linearly dependent on any sub-interval of $[T_{1},\infty)$. This contradicts Lemma 3.5. ∎
Lemma 3.7.
Suppose that $\mathcal{V}$ is a finite dimensional vector space of functions $f:[0,\infty)\to\mathbb{R}$ definable in a polynomially bounded o-minimal structure. Suppose that $\mathcal{V}$ does not contain non-zero functions, which are constantly zero eventually. Then there exist a basis $\{F_{1},...,F_{N}\}$ of $\mathcal{V}$, where
@@
We argue by induction. The claim is trivial for $N=1$. Now choose an arbitrary basis $\{H_{1},...,H_{N}\}$ of $\mathcal{V}$, and suppose without loss of generality that $\deg(H_{i})\leq\deg(H_{N})$ for all $i$. For each $i$ such that $\deg(H_{i})=\deg(H_{N})$, we have $\deg(H_{i}-c_{i}H_{N})<\deg(H_{N})$ for some $c_{i}\neq 0$, and we replace $H_{i}$ by $H_{i}-c_{i}H_{N}$. Now, the modified basis $\{H_{1},...,H_{N}\}$ is such that $\deg(H_{i})<\deg(H_{N}),\forall i<N$. The vector space $\mathcal{U}:=\text{Span}_{\mathbb{R}}\{H_{1},...,H_{N-1}\}$ is $(N-1)$-dimensional, and $\deg(f)<\deg(H_{N})$ for all $f\in\mathcal{U}$. By induction, there is a basis $\{F_{1},...,F_{N-1}\}$ for $\mathcal{U}$ such that $\deg(F_{i})<\deg(F_{i+1})$ for all $i$. In particular, $\{F_{1},...,F_{N-1},F_{N}\}$ with $F_{N}:=H_{N}$ is the required basis for $\mathcal{V}$. ∎
@@
If $\mathcal{V}$ is a finite dimensional vector spanned by definable functions, then the subset of functions which are eventually constantly zero forms a finite dimensional subspace $\mathcal{V}_{0}\leq\mathcal{V}$. This implies that there exists a uniform $T_{0}>0$ such that for all $f\in\mathcal{V}_{0}$, it holds that $f(t)=0,\forall t\geq T_{0}$. Thus, upon restricting the functions in $\mathcal{V}$ to $[T_{0},\infty)$, the subspace $\mathcal{V}_{0}$ is the trivial vector space.
We note the following fact:
  • A definable function $f:[0,\infty)\to\mathbb{R}$ either converges as $t\to\infty$ or diverges to $\infty$ or to $\infty$- ∞ as $t\to\infty,$ see [Geom_cat_and_o_min_str, Montonicity theorem, Section 4].
@@
Let $\mathcal{V}$ be a finite dimensional vector space of functions $f:[0,\infty)\to\mathbb{R}$ definable in a polynomially bounded o-minimal structure. Because of Remark 3.8, there is no loss in generality in assuming that the only eventually constantly zero function is the zero function. So we may choose a basis $\{F_{1}^{-},...,F_{n}^{-},F_{0}^{+},F_{1}^{+},...,F_{m}^{+}\}$ for $\mathcal{V}$ such that
$\deg(F_{1}^{-})<\cdots<\deg(F_{n}^{-})<0\leq\deg(F^{+}_{0})<\cdots<\deg(F_{m}^{+})$
, where $m,n\in\mathbb{N}\cup\{0\}$. Let $N=m+n+1$.
We will denote
$r_{i}=\deg(F^{+}_{i}),~{}-\kappa_{j}:=\deg(F^{-}_{j})$
where $r_{i},\kappa_{j}\geq 0$. For
$\underline{c}:=(c_{1}^{-},\ldots,c_{n}^{-},c_{0}^{+},\ldots,c^{+}_{m})\in\mathbb{R}^{N}$
we denote
(3.6) $f_{\underline{c}}:=c^{-}_{1}F^{-}_{1}+\cdots+c^{-}_{n}F^{-}_{n}+c^{+}_{0}F_{0}^{+}\cdots+c^{+}_{m}F^{+}_{m},$
(3.7) $f^{-}_{\underline{c}}:=c^{-}_{1}F^{-}_{1}+...+c^{-}_{n}F^{-}_{n},\text{ and }$
(3.8) $f^{+}_{\underline{c}}:=c^{+}_{0}F_{0}^{+}+...+c^{+}_{m}F^{+}_{m}.$
Let $\mathscr{F}\subseteq\mathcal{V}$ be a closed definable cone such that for all $0\neq f\in\mathscr{F}$ we have
$\lim_{t\to\infty}f(t)\neq 0.$
  • Let ~=$\tilde{\mathscr{F}}=\{\underline{c}\in\mathbb{R}^{n+m+1}:f_{\underline{c}}\in\mathscr{F}\}$, which is a closed cone in $\mathbb{R}^{m+n+1}$.
In addition, we assume (without loss of generality) that
$\lim_{t\to\infty}\frac{F^{+}_{i}(t)}{t^{r_{i}}}=1,~{}\lim_{t\to\infty}\frac{F^{-}_{j}(t)}{t^{-\kappa_{j}}}=1.$
We choose $T_{0}\geq 1$ such that the outcome of Lemma 3.6 holds. Fix $\delta\in(0,1)$, and assume for contradiction that $\mathscr{F}\smallsetminus\{0\}$ is not $\delta$-good on $[T_{0},\infty)$. Consider the following definable subset:
(3.9) $\mathcal{A}:=\left\{(s,\underline{c},a,l,\alpha):\begin{array}[]{c}s\geq 1,~{}\underline{c}\in\tilde{\mathscr{F}}\smallsetminus\{0\},~{}a>T_{0},~{}l>0,\\ a\leq\alpha\leq a+l-\delta l,~{}\frac{\|f_{\underline{c}}\|_{[a,a+l]}}{\|f_{\underline{c}}\|_{[\alpha,\alpha+\delta l]}}>s\\ \end{array}\right\}.$
By the assumption for contradiction, the projection of $\mathcal{A}$ to the first coordinate is $\mathbb{R}_{\geq 1}$. By the choice function theorem (cf. [Geom_cat_and_o_min_str, Section 4]), there exists a polynomially bounded definable curve

### v2

@@ 3. $(C,\alpha)$-good property
Let $\delta\in(0,1]$. A family $\tilde{\mathscr{F}}$ of real functions defined on $[T_{0},\infty)$ is called $\delta$-good if there exists a constant $M(\delta)$ depending only on $\delta$ such that
@@
for all $f\in\tilde{\mathscr{F}}$ and all bounded sub-intervals $I_{\delta}\subseteq I\subseteq[T_{0},\infty]$ satisfying $|I_{\delta}|=\delta|I|$.
Such inequalities are well-known in the literature as Remez-type inequalities; see [Remez_ineq].
We will prove the following result, from which Theorem 1.12 will be deduced. For the rest of this section, we fix a polynomially bounded o-minimal structure $\mathscr{S}_{\mathrm{pbd}}$.
@@
Let $\mathcal{V}$ be a finite-dimensional vector space of real functions defined on $[0,\infty)$ definable in $\mathscr{S}_{\mathrm{pbd}}$. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that
$\lim_{t\to\infty}f(t)\neq 0,\ \forall f\in\mathscr{F}\setminus\{0\}.$
Then, there exists $T_{0}>0$ such that the family of functions $\tilde{\mathscr{F}}$ of the nonzero functions in $\mathscr{F}$ restricted to $[T_{0},\infty)$ is $\delta$-good, for all $\delta\in(0,1]$.
Remark 3.3.
The $T_{0}$ in Theorem 3.2 can be chosen as a $T_{0}$, which is obtained by the outcome of Lemma 3.8.
@@
We first note the following corollary of Theorem 3.2.
@@
Let $\mathcal{V}$ be a finite-dimensional vector space of real functions on $[0,\infty)$ which are definable in $\mathscr{S}_{\mathrm{pbd}}$. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\setminus\{0\}$,
@@
Let $T_{0}\geq 1$ such that the outcome of Theorem 3.2 holds. Then, there exist $\mathrm{m},r>0$ such that:
(3.2) $\|f\|_{I}\leq\mathrm{m}x^{r}\|f\|_{I^{\prime}},$
for all $x\geq 1$, $f\in\mathscr{F}\setminus\{0\}$, and bounded intervals $I^{\prime}\subseteq I\subseteq[T_{0},\infty)$ such that $\frac{|I|}{|I^{\prime}|}\leq x$.
@@
Consider the definable set (see Lemma B.3):
$S:=\left\{(x,y):\forall I^{\prime}\subseteq I\subseteq[T_{0},\infty),~{}\frac{|I|}{|I^{\prime}|}\leq x,~{}\frac{\|f\|_{I}}{\|f\|_{I^{\prime}}}\leq y,~{}\forall 0\neq f\in\mathscr{F}\right\}.$
By Theorem 3.2, the projection of $S$ to the first coordinate includes $[1,\infty).$ Then, by the definable choice function theorem (see Theorem B.7), there exists a function $\phi:[1,\infty)\to\mathbb{R}^{2}$ definable in the same polynomially bounded o-minimal structure, such that
@@
Thus, for all $x>1$, $f\in\mathscr{F}\setminus\{0\}$, and intervals $I^{\prime}\subseteq I\subseteq[T_{0},\infty)$ such that $\frac{|I|}{|I^{\prime}|}\leq x$, it holds that
@@
Since $\phi$ is polynomially bounded, there exist $r>0$ and $T_{1}\geq 1$ such that $\frac{\phi(x)}{x^{r}}\leq c_{0}$ for all $x\geq T_{1}$, for some $c_{0}>0$. Finally, since $\frac{\phi(x)}{x^{r}}$ is bounded in $[1,T_{1}]$, the result follows. ∎
Our proof below is inspired by [Klein_Marg_flows_and_dioph, proof of Proposition 3.2].
Proof of Theorem 1.12 assuming Theorem 3.2.
We recall that the number of connected components of the family of definable sub-level sets
$\{t\in I:|f(t)|\leq\epsilon\},~{}\epsilon>0,~{}I\subseteq[T_{0},\infty),~{}f\in\mathscr{F}\setminus\{0\},$
is bounded uniformly, say by $K$, see Theorem B.6. Now fix $I\subseteq[T_{0},\infty),~{}\epsilon>0$ and an $f\in\mathscr{F}\setminus\{0\}$. Let
@@
$L\leq K|I^{\prime}|\text{, where }L:=|\{t\in I:|f(t)|\leq\epsilon\}|.$
Therefore, ${|I|}/{|I^{\prime}|}\leq{|I|}/{(L/K)}$. By Corollary 3.4, we get
@@
The space of polynomials of bounded degrees is the prototypical example of $\delta$-good functions. We record the following well-known result, as it will play a role in the proof of Theorem 3.2.
@@
Fix $n\in\mathbb{N}$. Then, for every $\delta\in(0,1]$, nested intervals $I_{\delta}\subseteq I$ with $|I_{\delta}|=\delta|I|>0$, and every nonzero polynomial $c_{0}+c_{1}x+\cdots+c_{n}x^{n}$, it holds that:
(3.3) $\frac{\|c_{0}+c_{1}x+\cdots+c_{n}x^{n}\|_{I}}{\|c_{0}+c_{1}x+\cdots+c_{n}x^{n}\|_{I_{\delta}}}\leq(n+1)\frac{n^{n}}{\delta^{n}}.$
We will be using the following convenient fact. Recall by the monotonicity theorem (Theorem B.5) that the derivative of a definable function $f(t)$ on an unbounded interval exists for all large $t$.
Lemma 3.6.
[Miller94, Proposition 3.1] Let $f:[0,\infty)\to\mathbb{R}$ be a function definable in $\mathscr{S}_{\mathrm{pbd}}$ such that $f(t)\sim t^{r}$ as $t\to\infty$, where $r\neq 0$. Then, $f^{\prime}(t)\sim rt^{r-1},\text{ as }t\to\infty$.
Lemma 3.7.
Let $F_{i}:[0,\infty)\to\mathbb{R}$ be functions definable in $\mathscr{S}_{\mathrm{pbd}}$, for $i=0,1,\ldots,N$, and assume that none of the $F_{i}$ are eventually zero. Suppose that the degrees $\deg(F_{0}),\ldots,\deg(F_{N})$ are all distinct. Then, there exists a $T_{0}>0$ such that for any interval $I\subseteq[T_{0},\infty)$ of positive length, it holds that
$\|c_{0}F_{0}+c_{1}F_{1}+...+c_{N}F_{N}\|_{I}>0$
for all $(c_{0},c_{1},\ldots,c_{N})\neq 0$.
@@
By o-minimality (Theorem B.6), the definable set $F_{0}^{-1}(0)$ has finitely many connected components. Since $F_{0}$ is not eventually zero, for some $T_{0}>0$, $F_{0}(t)\neq 0$ for all $t\geq T_{0}$. Thus for all $I\subseteq[T_{0},\infty)$ we have
$\left\|c_{0}F_{0}+c_{1}F_{1}+\cdots+c_{N}F_{N}\right\|_{I}>0\iff\left\|c_{0}+c_{1}\frac{F_{1}}{F_{0}}+\cdots+c_{N}\frac{F_{N}}{F_{0}}\right\|_{I}>0.$
Since the degrees of $F_{0},F_{1},...,F_{N}$ are distinct, the degrees of
$1,F_{1}/F_{0},...,F_{N}/F_{0},$
are distinct. Thus, we are reduced to proving the statement under the assumption that $F_{0}(t)=1$ for all $t\geq T_{0}$.
Pick $T_{1}\geq T_{0}$ such that the functions $F_{1},\ldots,F_{N}$ are differentiable in $[T_{1},\infty)$, see Theorem B.5. By Lemma 3.6, $\deg(F_{i}^{\prime})=\deg(F_{i})-1$ and $F_{i}^{\prime}$ is not eventually zero for $1\leq i\leq N$. So, the degrees of $F_{1}^{\prime},F_{2}^{\prime},\cdots,F_{N}^{\prime}$ are all distinct. Now, for any interval $I\subseteq[T_{1},\infty)$ of positive length,
if $\text{if }\|c_{0}+c_{1}F_{1}+...+c_{N}F_{N}\|_{I}=0\text{, then }\|c_{1}F_{1}^{\prime}+...+c_{N}F_{N}^{\prime}\|_{I}=0.$
In view of the above observations, the result is straightforward to deduce by induction. ∎
Next, for definable functions $F_{0}(t),F_{1}(t),...,F_{N}(t)$, we consider the Wronskian matrix $W(F_{0},...,F_{N})(t)$, which is defined to be the $(N+1)\times(N+1)$ matrix whose $k$-th row is $(F^{(k)}_{0}(t),F_{1}^{(k)}(t),\ldots,F_{N}^{(k)}(t))$ for $0\leq k\leq N$. The Wronskian is well defined for all $t$ large enough due to Theorem B.5.
@@
Let the notation be as in Lemma 3.7. Then, there exists a $T_{0}>0$ such that the functions $F_{0},...,F_{N}$ are $(N+1)$-time continuously differentiable and $W(F_{0},...,F_{N})(t)$ is non-singular for all $t\geq T_{0}$.
@@
Since $\det(W(F_{0},...,F_{N})(t))$ is a definable function, either
$|\det(W(F_{0},...,F_{N})(t))|>0,$
for all large $t$ or $\det(W(F_{0},...,F_{N})(t))=0$ for all large $t$. We will prove the result by induction. Since $F_{0}$ is not eventually zero, the result follows for $N=0$. Now suppose $N\in\mathbb{N}$ and $T_{0}>0$ such that $\det(W(F_{0},...,F_{N-1})(t))\neq 0$ for all $t\geq T_{0}$. Now, suppose $\det(W(F_{0},...,F_{N})(t))=0$ for all $t\geq T_{1}\geq T_{0}$. Then, by a classical result of Bôcher [Bochner_certain_cases_Wronksian_implies_dep, Theorem II], $F_{0},\ldots,F_{N}$ are linearly dependent on $[T_{1},\infty)$. This contradicts Lemma 3.7. ∎
@@
Let $\mathcal{V}$ be a real finite-dimensional vector space spanned by functions definable in $\mathscr{S}_{\mathrm{pbd}}$, and suppose that the zero function is the only function in $\mathcal{V}$ which is eventually constantly zero. Then, there exists a basis $\{F_{1},...,F_{N}\}$ of $\mathcal{V}$ such that:
@@
We argue by induction. The claim is trivial for $N=1$. Now choose an arbitrary basis $\{H_{1},...,H_{N}\}$ of $\mathcal{V}$, and suppose without loss of generality that $\deg(H_{i})\leq\deg(H_{N})$ for all $i$. Moreover, we claim that there is no loss of generality in assuming that $\deg(H_{i})<\deg(H_{N})$, for all $i<N$. Indeed, for each $i<N$ such that $\deg(H_{i})=\deg(H_{N})$, we can find $c_{i}\neq 0$ such that $\deg(H_{i}-c_{i}H_{N})<\deg(H_{N})$. By replacing $H_{i}$ with $H_{i}-c_{i}H_{N}$, we again get a basis for $\mathcal{V}$. Note that the vector space $\mathcal{U}:=\text{Span}_{\mathbb{R}}\{H_{1},...,H_{N-1}\}$ is $(N-1)$-dimensional, and $\deg(f)<\deg(H_{N})$ for all $f\in\mathcal{U}$. By induction, there is a basis $\{F_{1},...,F_{N-1}\}$ for $\mathcal{U}$ such that $\deg(F_{i})<\deg(F_{i+1})$ for all $i$. Then, $\{F_{1},...,F_{N-1},F_{N}\}$ with $F_{N}:=H_{N}$ is the required basis for $\mathcal{V}$. ∎
@@
If $\mathcal{V}$ is a finite-dimensional vector spanned by definable functions, then the subset of functions which are eventually constantly zero forms a finite-dimensional subspace $\mathcal{V}_{0}\subseteq\mathcal{V}$. Then, there exists a uniform $T_{0}>0$ such that for all $f\in\mathcal{V}_{0}$, it holds that $f(t)=0$, for all $t\geq T_{0}$. Thus, by restricting the functions in $\mathcal{V}$ to $[T_{0},\infty)$, we may assume $\mathcal{V}_{0}=\{0\}$.
@@
Let $\mathcal{V}$ be a finite-dimensional vector space of functions $f:[0,\infty)\to\mathbb{R}$ definable in a polynomially bounded o-minimal structure. By Remark 3.10, there is no loss in generality in assuming that only the zero function is eventually constantly zero. Using Lemma 3.9, we may further assume, by possibly enlarging $\mathcal{V}$, that $\mathcal{V}$ has a basis of the form $\{F_{1}^{-},...,F_{n}^{-},F_{0}^{+},F_{1}^{+},...,F_{m}^{+}\}$ such that
$\deg(F_{1}^{-})<\cdots<\deg(F_{n}^{-})<0\leq\deg(F^{+}_{0})<\cdots<\deg(F_{m}^{+}),$
where $m,n\in\mathbb{N}$. Let $N=m+n+1$. We will denote
(3.5) $r_{i}=\deg(F^{+}_{i}),\quad-\kappa_{j}:=\deg(F^{-}_{j}),$
where $r_{i},\kappa_{j}\geq 0$. In addition, we assume for convenience that
$\lim_{t\to\infty}F^{+}_{i}(t)/t^{r_{i}}=1\text{ and }\lim_{t\to\infty}F^{-}_{j}(t)/t^{-\kappa_{j}}=1.$
For any $\underline{c}:=(c_{1}^{-},\ldots,c_{n}^{-},c_{0}^{+},\ldots,c^{+}_{m})\in\mathbb{R}^{N}$, we denote
$f^{-}_{\underline{c}}:=c^{-}_{1}F^{-}_{1}+\cdots+c^{-}_{n}F^{-}_{n},\ f^{+}_{\underline{c}}:=c^{+}_{0}F_{0}^{+}+\cdots+c^{+}_{m}F^{+}_{m}\text{, and }f_{\underline{c}}=f^{-}_{\underline{c}}+f^{+}_{\underline{c}}.$
Let $\mathscr{F}\subseteq\mathcal{V}$ be a closed definable cone such that
$\lim_{t\to\infty}f(t)\neq 0,\,\forall f\in\mathscr{F}\setminus\{0\}.$
Let ~:=$\tilde{\mathscr{F}}:=\{\underline{c}\in\mathbb{R}^{N}:f_{\underline{c}}\in\mathscr{F}\}\cong\mathscr{F}$. Then $\tilde{\mathscr{F}}$ is a closed cone in $\mathbb{R}^{N}$.
Fix a $T_{0}\geq 1$ for which the outcome of Lemma 3.8 holds, and suppose for contradiction that there exists a $\delta\in(0,1)$ such that $\mathscr{F}\setminus\{0\}$ is not $\delta$-good on $[T_{0},\infty)$. Consider the following definable subset (see Lemma B.3):
(3.6) $\mathcal{A}:=\left\{(s,\underline{c},a,l,\alpha):\begin{array}[]{c}s\geq 1,~{}\underline{c}\in\tilde{\mathscr{F}}\setminus\{0\},~{}a>T_{0},~{}l>0,\\ a\leq\alpha\leq a+l-\delta l,~{}\frac{\|f_{\underline{c}}\|_{[a,a+l]}}{\|f_{\underline{c}}\|_{[\alpha,\alpha+\delta l]}}>s\\ \end{array}\right\}.$
By the assumption for contradiction, the projection of $\mathcal{A}$ to the first coordinate equals $[1,\infty)$. By the choice function theorem (Theorem B.7), there exists a polynomially bounded definable curve